# SiriusBI: Building End-to-End Business Intelligence Enhanced by Large Language Models

### v1

@@ 2.3. Knowledge Management
Knowledge Management (KM) involves systematic processes within an organization aimed at optimizing the use of knowledge. In most BI and ChatBI systems, KM is structured into three modules: knowledge processing, knowledge storage, and knowledge retrieval.
The critical formula for knowledge retrieval is:
where represents the structured knowledge from retrieval, represents the user’s query, denotes the stored knowledge, and is the function that facilitates the retrieval of knowledge from storage.
2.4. Data Insight
Data insights is a fundamental component of BI and ChatBI systems, designed to transform raw data into actionable insights through a series of structured processes: data aggregation, data analysis, insight generation, and visualization.
Data Aggregation: Combine data from multiple sources into a comprehensive dataset.
Data Analysis: Apply statistical and machine learning techniques to identify patterns and trends within the aggregated data.
Insight Generation: Translate the results of data analysis into meaningful insights that can inform business decisions.
Visualization: Present insights through intuitive visual formats to enhance understanding and facilitate decision-making.
3. System design
In this section, we introduce the overall architecture of SiriusBI. Compared with the current ChatBI systems (Lian et al., 2024), SiriusBI provides a more comprehensive workflow, especially in the knowledge management and data insight.
3.1. Overview
The overall framework of SiriusBI is illustrated in Figure 2 and consists of five key modules: (1) The Knowledge Management subsystem which stores, processes, and retrieves pertinent database knowledge, domain-specific expertise, and demonstrations (i.e., examples used for context learning in LLMs) that support the functionality of the other modules. (2) The Dialogue Analysis module, which accepts user queries and handles the interaction through multiple-round dialogues. (3) The Intention Querying module, which clarifies the user’s intent by generating follow-up questions based on the dialogue history. (4) The SQL Generation module, which converts the fully understood user intent into SQL queries. (5) The Data Insight module, which automatically prepares data, plans tasks, attributes issues, and predicts trends based on the data retrieved through NL2SQL queries.
3.2. Workflow
The workflow of SiriusBI is detailed as follows: When a user inputs a question in a BI scenario, for example, ”Company A’s year-to-date consumption compared year-over-year?” (), in the first stage, the Dialogue Analysis module evaluates whether the semantic content of the user’s input is complete, i.e., whether it includes both ”dimensions” and ”metrics”. If the input is determined to be semantically complete, it proceeds to the second stage, the Intention Querying module. However, if the user’s input is missing either dimensions or metrics, the module uses the conversation history to complete the semantics of the current question before proceeding to the next stage. For instance, if the user subsequently asks, ”How about Company B?” (), the Dialogue Analysis module will complete the question to ”Company B’s year-to-date expenditure compared year-over-year?” ().
In the second stage, the Intention Querying module recalls relevant information from the knowledge base based on the input question. Notably, to incorporate domain knowledge, improve system generalization, and enhance data insights, we have developed a specialized Knowledge Management module, which is elaborated in 4.1. After knowledge retrieval, we recall relevant fields from the database. If there is still ambiguity or confusion among the recalled fields, the Intention Querying module prompts the user for clarification to ensure accurate input for the SQL generation stage, thereby minimizing performance loss due to error propagation in the pipeline.
In the third stage, the SQL generation module receives the semantically completed user query from the previous stage and selects the appropriate tables for each query. Typically, this involves matching the top-k tables using database table information. To improve the adaptability of SiriusBI to various scenarios and requirements, we propose two optional generation methods: 1) An one-step generation method, where we construct specific datasets via an automated process for and perform Supervised Fine-Tuning (SFT) on tasks that require strong domain expertise but lower generalization; 2) A two-step SQL generation method, where we structure user queries via knowledge retrieval and then translate inputs into SQL using a fixed LLM. This method ensures fast deployment of SQL generation on unseen scenarios. By parallel use of these two generation methods, we maximize the accuracy of the generated SQL queries, ensuring reliable and relevant results for users.
In the fourth stage, the Data Insight module receives the SQL queries and their corresponding results from the third stage. Based on the user’s additional insight requirements, the module performs automatic data preparation, complex task planning, insight tool execution, and finally generates corresponding reports.
In the following section, we provide more in-depth exploration of the design and functionalities of each module.
4. System description
In this section, we explore various components of SiriusBI in detail. We begin with the Knowledge Management module in Section 4.1, which serves as a critical foundation that provides essential knowledge support for other parts of the workflow. Next, Section 4.2 focuses on Multi-round Dialogue with Querying, which consists of both the Dialogue Analysis and Intention querying module. In Section 4.3, we discuss SQL Generation, highlighting both the one-step and two-step SQL generation schemes. Finally, we introduce the Data Insight module in Section 4.4.
4.1. Knowledge Management
In industrial scenarios, the application of Business Intelligence (BI) often requires a substantial amount of domain-specific knowledge, including table metadata, business knowledge, and LLMs chat history, etc. To address the complex knowledge requirements, SiriusBI establishes the Knowledge Management module by constructing a multidimensional knowledge graph.
This knowledge graph comprises several key dimensions: 1) metadata information, which includes table schema, column values; 2) business knowledge, which includes domain-specific fields, term definitions, and relevant business context; 3) demonstration information utilized for prompting; 4) conversation data. Together, these dimensions establish a robust foundation for the knowledge management module.
The module also supports DSL (Domain-Specific Language) data generation, which includes functionalities such as query rewriting and DSL prompting. Furthermore, it facilitates API data retrieval and provides essential metadata and business knowledge, thereby supporting comprehensive data insights. This versatility makes the knowledge management module a vital component in various data-driven applications.
@@ 4.1.1. Knowledge Storage
For knowledge storage, we use EasyGraph (Gao et al., 2023) to construct a multidimensional knowledge graph. Figure 3 shows the graph stucture for storing knowledge related to data tables. Each node is a data structure that contains node types, descriptions, embeddings, and other information. The node types in this graph include databases, tables, columns, (column) values, UDFs (complex metric calculation functions such as month-over-month), terms (explanations of professional terms), and aliases (alternative names for other nodes). For columns, the node has specific attributes like whether it is a dimension or measure. The relationships between these nodes are represented as edges, forming a comprehensive knowledge storage structure used in SiriusBI. By defining graph structures and enhancing query efficiency through advanced graph algorithms, we not only organize knowledge effectively but also facilitate the rapid retrieval of relevant information during queries.
4.1.2. Knowledge Processing
Before storing the knowledge, it’s essential to process various types of data to ensure effective storage and management. We provide standard guidelines for the business teams to follow. To create a new node, domain experts first manually extract the node attributes based on specific business scenarios. Then, for knowledge with missing attributes, we complete it through metadata enhancement (Choudhury et al., 2023). This enhancement occurs on two levels: 1) automatic context completion using AI techniques, and 2) backfilling missing metadata through data engineering methods. Together, these processes ensure that the knowledge graph is ready for effective utilization.
4.1.3. Knowledge Retrieval
The knowledge retrieval process is divided into two main phases: coarse retrieval and fine retrieval. Figure 4 shows a brief overview of the retrieval process.
Coarse Retrieval. Initially, we use Elasticsearch (Elasticsearch, 2018) for coarse retrieval without segmenting the original query. Specifically, we deploy a hybrid search on both the lexical and semantic levels (Wang et al., 2024). To achieve a high recall rate, we allow for a larger recall set, temporarily prioritizing recall over precision. For each recalled node, if it is an alias or a value, we trace back to at least the column, term, or UDF level. For example, for SQL generation, this process yields all potentially relevant column names for fine retrieval.
Fine Retrieval. Based on the results from coarse retrieval, we perform a secondary search on the sub-graph obtained from the coarse retrieval by selecting the top-k nodes ranked by a reranking model (Chen et al., 2024b), which significantly reduces the number of candidate nodes. In addition to reranking, for SQL generation, we apply a LLM to filter out columns if it is not relevant to the user query.
@@ 4.2. Multi-Round Dialogue with Querying
Multi-Round Dialogue with Querying (MRD-Q) is a critical component in SiriusBI that receives and processes user input. It primarily supports multi-round dialogue (MRD), intent recognition, and clarification aimed at the user. This reduces the usage barrier, enhances the user experience, and minimizes the impact of errors caused by incorrect intent recognition. MRD-Q consists of two modules: dialogue analysis and intention querying. A detailed introduction is as follows:
@@ 4.2.1. Dialogue Analysis
In the industrial sector, BI has a wide range of applications aimed at enhancing user experience in data retrieval, data analysis, and data science. In this context, the multi-round dialogue feature in ChatBI plays a crucial role. However, most current NL2SQL research (Pourreza and Rafiei, 2023; Wang et al., 2023; Sun et al., 2023; Caferoğlu and Ulusoy, 2024; Li et al., 2024) is designed based on single-round dialogues, which creates a significant gap between this approach and real-world application scenarios. Firstly, single-round NL2SQL tasks require highly precise user input, meaning that even minor errors in the input may prevent the generation of correct SQL statements for subsequent data processing tasks. This high requirement not only severely diminishes user experience but also raises the barrier to entry for users who may not be familiar with SQL syntax. Secondly, data tasks, including data retrieval, are not isolated processes. Real-world data analysis or data science tasks often involve complex and detailed work, which inevitably requires extensive data retrieval efforts and even more sophisticated BI functionalities. The single-round dialogue approach for data retrieval is evidently inadequate to meet the demands of these real-world application scenarios.
The multi-round dialogue feature in SiriusBI is built upon the basic principles of dialogue analysis implementation in ChatBI(Lian et al., 2024). It determines the input for the SQL Generation module by assessing whether the user’s current input or historical conversation records contain indicators (columns) and dimensions. Specifically, the system first determines whether the user’s current input contains the two key pieces of information: indicators and dimensions. If both are found, the current input is considered semantically complete, and the workflow proceeds to the next module. If either is missing, the system supplements the current input’s semantics by selecting the most recent historical input that includes indicators and dimensions, and then moves to the next module.
Notably, unlike ChatBI (Lian et al., 2024), SiriusBI does not use traditional small NLP model fine-tuning to determine whether a statement contains indicators and dimensions or to select relevent historical conversations. Instead, we leverage the robust semantic understanding capabilities of LLMs along with the domain-specific knowledge in our knowledge base. This approach not only facilitates the system’s transition between different vertical domains but also enhances SiriusBI’s adaptability across various fields, which leads to a reduction on time and data construction costs associated with model training.
@@ 4.2.2. Intention Querying
To address the interaction limitations, we have added an intent clarification feature to the basic dialogue analysis module, resulting in a comprehensive MRD-Q component. The intention querying module is a critical supplement to dialogue analysis and includes the following three steps:
Intent Classification. When processing information from dialogue analysis, the first step is to reclassify the user’s true intentions. In multi-turn dialogues in real-world scenarios, certain metrics or dimensions may be missing, and relying solely on the dialogue history cannot complete the information. At this point, the results of intent classification are needed to decide whether to ask the user follow-up questions. We use the same LLM for semantic understanding as in Section 4.2.1 and categorize queries into three types. Denote the user input as , and the result of the intent classification as . We have,
When equals to 1, it indicates that, despite supplementing with historical inputs in Section 4.2.1, the user query still lacks critical information. Therefore, to ensure the successful execution of the task, we must either automatically retrieve essential information from the knowledge base or request additional details from the user.
Knowledge Retrieval. We utilize the knowledge retrieval function described in Section 4.1.3, using a hybrid approach of Embedding and ElasticSearch to retrieve demonstrations and domain knowledge. The demonstrations primarily serve as query clarification prompts. We have manually constructed a prompt library of common query clarification forms and stored it along with the necessary knowledge for SiriusBI in the knowledge management module.
Intent Clarification. Based on the results of intent recognition and knowledge retrieval, SiriusBI integrates and comprehensively analyzes these key pieces of information during the clarification phase. Specifically,
When is 0, the clarification phase will politely decline and guide the user to ask data analysis-related questions.
When is 1, the input is likely not reflected in the user’s historical information. In this case, the clarification phase will rely on the retrieved information from the knowledge base to enhance the indicators and dimensions. It will then ask the user if the supplemented parts meet their expectations. If not, the user is prompted to provide the relevant indicators and dimensions in the next round of dialogue. This iterative process helps clarify the user’s true intent through follow-up questions, i.e., the true dimensions and indicators of the query.
When is 2, no further clarification is needed, and the process directly moves to the next module in the SiriusBI pipeline.
By integrating the intent clarification feature into the dialogue analysis, we have developed the MRD-Q component. Through professionally designed clarification prompts, we can understand the user’s true intent more accurately. This approach lowers the user’s entry barrier. Even if the user provides arbitrary answers, SiriusBI can clarify the user’s correct intent through guided follow-up questions. Additionally, it helps to avoid confusion between similar dimensions and indicators, thereby preventing error propagation throughout the entire BI system.
@@ 4.3. SQL Generation
NL2SQL is the core task within the entire BI system, and the performance bottleneck continues to pose challenges in the industry. In this subsection, we will focus on the implementation approach of the SQL Generation module in SiriusBI, including table selection, as well as the SQL generation methods. To address the high cost of model migration across different professional domains, we have also proposed a novel two-step paradigm for SQL generation.
@@ 4.3.1. Table Selection
Before generating SQL queries, it is necessary to select the appropriate tables from the database based on the user’s input, especially retrieving the relevant table information mentioned in the user’s query. As shown in Figure 5, the implementation of table selection is divided into two parts: coarse ranking and reranking. In the coarse ranking stage, the number of candidate tables is reduced to a manageable number. In the reranking stage, these candidates are re-ranked to identify the most relevant ones for SQL generation.
The coarse ranking process consists of two parts: online and offline. In the offline part, each data table has a pre-prepared embedding. We use the StarRocks (SR) (team, [n.d.]) vector database to store the embeddings of tables to ensure data security. SR supports millisecond-level latency retrieval and allows for filtering by specified tags, making it a robust solution for efficient table selection. In the online part, we extract multiple keywords from the query through a LLM and then perform individual retrieval using each keyword. Compared to directly embedding the user’s input, this approach eliminates redundant information, reduces interference from irrelevant information, and increases the recall rate of target tables.
During reranking, we use different dimensional scoring methods for each table based on the hundreds of tables recalled in the coarse ranking stage. The overall calculation formula is as follows:
@@
We utilize TF-IDF (Joachims, 1997) as the vectorization method to compute similarity for several reasons: 1) The keywords are relatively short; 2) The calculation needs to be performed in real-time; 3) Gensim-based TF-IDF allows for rapid computation of similarity in batches; 4) Avoid the situation where the vectorized results fail to give proper attention to some keywords in the query when the input is excessively long.

### v2

@@ 3.2. Knowledge Management
The application of BI often requires substantial amount of domain-specific knowledge. To address the knowledge requirements, SiriusBI constructs and maintains a comprehensive knowledge base within the knowledge management module. This knowledge base is designed to support efficient data retrieval, enhance decision-making processes, and facilitate seamless integration with BI tools.
The knowledge base primarily constitutes four types of data: (1) Database Metadata: This includes essential information such as table names, column types, column names, and other details of the database schema. (2) Domain-Specific Knowledge: This encompasses explanations of certain fields, term definitions, and relevant business context. It provides the necessary background information to interpret and utilize the data effectively within the specific domain. (3) Task-Specific Materials: These are groups of prompts, demonstrations and other resources designed for different tasks. They are tailored to guide users in performing specific operations or analyses within the BI system. (4) Historical Conversations: This includes logs of past interactions and queries, which can be used to complete semantic for multi-round dialogue analysis.
@@ 3.2.1. Knowledge Storage
We observe that certain types of knowledge entries exhibit dependencies, while others remain relatively independent. In our application scenario, the relationships between database metadata and relevant domain-specific knowledge can be represented as a directed acyclic graph as illustrated in Figure 4. To ensure comprehensive retrieval of all related knowledge, it is essential to explicitly capture and represent these relationships.
While graph databases (Gao et al., 2023) offer a natural solution for storing such relational knowledge, we opt against their use due to their inherent complexity and overhead, which exceed the requirements of our system. Instead, we adopt a lightweight approach by augmenting knowledge entries with two dedicated fields: Anc. Label, which specifies the label of the ancestor entry, and Anc. Name, which encodes the concrete value of the ancestor entry. This design, as depicted in Figure 5 (right table), preserves the relational semantics necessary for knowledge retrieval while avoiding the operational and computational complexities associated with native graph storage. This approach strikes a balance between functionality and efficiency, aligning with the practical needs of our system.
For independent knowledge entries, we simply store a description of the entry and, if applicable, a demonstration or example. This straightforward approach is sufficient for entries that do not have dependencies or relationships with other knowledge items, as shown in Figure 5 (left table).
3.2.2. Knowledge Extraction
The objective of the knowledge extraction process in SiriusBI is to efficiently and accurately extract relevant information from the knowledge base. We employ different knowledge extraction strategies based on the storage strategy described earlier. All queries must undergo at least one round of knowledge retrieval, which is divided into two main phases: coarse retrieval and fine retrieval. This two-phase approach ensures a balance between recall and precision, enabling the system to first identify a broad set of potentially relevant knowledge and then refine the results to meet the specific needs of the task at hand.
In the coarse retrieval phase, the primary goal is to maximize recall, ensuring that a comprehensive set of potentially relevant knowledge entries is retrieved. This is particularly important for tasks such as SQL generation, where the system needs to identify all possible column names or metadata that might be relevant to the user’s query. To achieve this, we employ a mixed-metric search strategy that operates on both lexical and semantic levels (Wang et al., 2024). Lexical search captures exact matches or close variants of the query terms, while semantic search leverages embeddings to identify conceptually related knowledge entries, even if they do not share exact lexical overlap. For this phase, we utilize Elasticsearch (Elasticsearch, 2018), a widely adopted search engine known for its scalability and efficiency. The query is processed without segmentation, allowing for a broader initial recall set.
For knowledge entries that exhibit hierarchical relationships (e.g., those with Anc. Name fields), we implement a multi-round retrieval strategy inspired by multi-hop graph traversal. Specifically, for each entry retrieved in the initial search, if its Anc. Name field is non-empty, we recursively retrieve the ancestor entries. This process continues iteratively until no further ancestor entries are found. This approach ensures that the retrieval process captures not only the directly relevant knowledge but also the contextual and hierarchical relationships that are critical for tasks such as SQL generation or business rule interpretation.
To illustrate this process, consider the right table in Figure 4. The retrieval begins with the entry “Campus Hiring”, whose Anc. Name field contains the ancestor entry “Campus Recruitment”. The system then retrieves “Campus Recruitment”, whose Anc. Name field includes “Recruitment Channels”. Next, “Recruitment Channels” is retrieved, which further links to “Personnel Table”. Finally, “Personnel Table” is retrieved, whose Anc. Name field points to “Company HR Database”. Since “Company HR Database” is not present in the knowledge base, the retrieval process terminates. This example demonstrates how the multi-hop traversal strategy effectively captures a chain of related knowledge entries, ensuring comprehensive retrieval of contextually relevant information.
The coarse retrieval phase typically yields a large set of candidate knowledge entries. To refine these results, we perform a fine retrieval phase, which focuses on improving precision by re-ranking and filtering the candidate set. We employ a state-of-the-art re-ranking model, BGE (Chen et al., 2024b), to rank the candidate nodes based on their relevance to the query. For tasks such as SQL generation, we further refine the results by leveraging a LLM to filter out irrelevant columns. The LLM evaluates the semantic relevance of each column to the user query and removes those that do not contribute to the intended SQL query. This step ensures that the final set of retrieved knowledge is both concise and highly relevant to the task.
By combining coarse retrieval, relationship-aware multi-hop traversal, re-ranking, and LLM-based filtering, SiriusBI achieves a robust and efficient knowledge retrieval process. This approach not only ensures high recall and precision but also adapts to the complex dependencies and relationships inherent in domain-specific knowledge, making it well-suited for a wide range of BI applications.
@@ 3.3. Multi-Round Dialogue with Querying
Multi-Round Dialogue with Querying (MRD-Q) is a critical component in SiriusBI that receives and process user inputs for data integrity. It is designed to support MRD through automatic semantic completion and user-involved intent querying. A detailed introduction is presented in the following sections.
@@ 3.3.1. Dialogue Analysis
The multi-round dialogue analysis mechanism in SiriusBI is built upon the basic dialogue analysis module in previous work(Lian et al., 2024). It first determines the semantic integrity of the user query by assessing whether the current input or historical conversations contain metrics and dimensions. In the context of business intelligence, dimensions refer to attributes that provide granularity level information of the user query, such as time periods or geographical locations. For example, “last quarter” is a time dimension and “in New York” is of location dimension. Metrics, on the other hand, are quantitative measurements, such as the number of sales, which might be represented by a column like Total_Sales. When a query includes both dimensions and metrics, such as “the number of sales in New York last month”, it is possible to construct a corresponding SQL query. Conversely, if a query lacks either a dimension or a metric, such as “How about Los Angeles?”, it becomes challenging to generate a meaningful SQL query. Although this approach is somewhat intuitive, it effectively addresses most dialogue patterns encountered in BI scenarios.
If both are found, the current input is considered of semantic integrity, and the workflow proceeds to the next module. If either is missing, the system supplements the current input’s semantics by selecting the most recent historical input that includes metrics and dimensions, and then moves to the next step.
Unlike previous work, SiriusBI does not fine-tune a PLM (Pre-trained Language Model, e.g., (Lian et al., 2024) adopts ERNIE (Sun et al., 2019)) to determine the semantic integrity or to select relevant historical conversations. PLMs, such as BERT (Devlin et al., 2019), GPT-2 (Radford et al., 2019), and ERNIE (Sun et al., 2019), are large-scale neural networks pre-trained on vast amounts of text data to capture general language representations. While they have shown remarkable performance in various natural language processing tasks, fine-tuning them for specific applications often demands substantial labeled data and computational resources. Additionally, fine-tuned models may not generalize well across different domains, limiting their transferability and effectiveness in diverse BI scenarios. Instead, we leverage the robust reasoning capability of LLMs to figure out the existence of “metric” and “dimension” and the historical conversations.
@@ 3.3.2. Intention Querying
For multi-round dialogues in real-world scenarios, relying solely on the dialogue history is often insufficient to complete the missing information due to the inherent complexity. To address these challenges, we enhance the basic dialogue analysis module by introducing an intent querying feature, resulting in a robust MRD-Q component. The intention querying process consists of the following steps:
Intent Classification. We use the same LLM for dialogue analysis in Section 3.3.1 and further categorize the expanded queries into three types. Denote the user input as , and the result of the intent classification as . Formally, we classify the query into three categories, as indicated by Eqn. (1).
When , the query is identified as out of the BI scope. In this case, the system will politely decline the request and guide the user to ask data analysis-related questions. The dialogue analysis process will then restart. When , it indicates that, even after supplementing with historical inputs, the user query still lacks key information. The system will proactively ask the user to provide the missing information and then restart the analysis process. When , the query is complete. The system will proceed to the next step of the pipeline.
Intent Clarification. Although the queries entering this stage are semantically complete, additional processing is required to ensure the generated SQL accurately reflects the user’s intent due to real-world complexities. To address this, the module first retrieves domain-specific knowledge from the knowledge base based on the user’s query. For instance, the triplet associated with the term “YTD” is defined as: (“YTD”, “Year To Date or Year To Day, referring to the period from the beginning of the current year to the present date.”, “What is the YTD revenue of the mini-program?”). Such knowledge enhances the LLM’s understanding of domain-specific terminology. Furthermore, the module retrieves task-specific materials, including prompt templates and demonstration examples, to guide the LLM in intent clarification. Specifically, the LLM either returns a semantically refined query or interactively presents users with disambiguation options to ensure input precision.
In summary, we develop the MRD-Q component, which supplements the heuristic dialogue analysis by the intent clarification capability of the LLM. Even if the user provides arbitrary answers, SiriusBI can clarify the user’s true intent through guided follow-up questions, therefore improving the successful rate of the task execution.
@@ 3.4. SQL Generation
NL2SQL is the core task within the entire BI system, and the performance bottleneck continues to pose challenges in the industry. In this subsection, we will focus on the implementation approach of the SQL Generation module in SiriusBI, including table selection, as well as the SQL generation methods.
@@ 3.4.1. Table Selection
Before generating SQL queries, it is necessary to select the appropriate tables from numerous options in an industrial database based on the user’s input. This process relies on retrieving the relevant table information mentioned in the user’s query. The implementation of table selection is divided into two parts: coarse ranking and re-ranking. During the coarse-ranking stage, the number of candidate tables is typically reduced by a factor of ten or more to maintain fewer than 100 candidates, thereby ensuring computational efficiency in the subsequent fine-ranking phase. In the re-ranking stage, these candidates are re-ranked to identify the most relevant ones for SQL generation.
The coarse ranking process consists of two parts: online and offline. In the offline part, each data table is subjected to a private embedding model and transformed into an embedding vector. We use the StarRocks (SR) (team, 2023b) vector database to store the embeddings of tables to ensure data security and the embeddings are extracted using the general-purpose pre-trained model m3e-large (Chen et al., 2024c), ensuring high-quality feature representations and domain adaptability. SR supports millisecond-level latency retrieval and allows for filtering by specified tags, making it a robust solution for efficient table selection. In the online part, we extract multiple keywords from the query through a LLM and then perform individual retrieval using each keyword. Compared to directly embedding the user’s input, this approach eliminates redundant information, reduces interference from irrelevant information, and increases the recall rate of target tables.
During re-ranking, we use a scoring function balancing multiple metrics for each table based on the hundreds of tables recalled in the coarse ranking stage. The overall calculation formula is as follows:
@@
While TF-IDF (Joachims, 1997) remains effective for real-time matching on short text queries, we perform coarse filtering by restricting candidate tables to those with dynamically adjusted per query complexity.
3.4.2. Dynamic Strategy Switching Driven by Data Conditions
In practical BI applications, the data conditions of different clients vary significantly: some clients have abundant labeled data, while others can only provide a limited number of labeled samples due to frequent business changes or data privacy restrictions. Similarly, existing SQL generation methods are often tailored to specific scenarios or domains, making migration between private scenarios costly in terms of both deployment effort and data labeling. To address this challenge, we propose a data-condition-driven dynamic strategy switching mechanism. This mechanism adaptively selects either a one-stage (high-precision fine-tuning) or two-stage (semantic inter- mediate representation) generation strategy based on two key factors: the volume of labeled data and the semantic similarity between domains. By doing so, it effectively balances accuracy and cost-efficiency.
Formally, we define the number of available labeled Query-SQL pairs as the Labeled Data Volume(). We also quantify the semantic similarity between the target domain and an existing labeled domain (referred to as the source domain) as Domain Similarity()This similarity is computed via the cosine similarity between the embeddings of schema keywords provided by the business side, as follows:
where and represent the sets of keywords for the target domain and the source domain, respectively (e.g., in the financial domain: ”revenue”, ”quarterly report”). is Sentence-BERT (Reimers and Gurevych, 2019) encoding function.
Based on the aforementioned features, we use rule-based criteria to determine the SQL generation strategy. Specifically:
where and are hyperparameters that are typically specified based on the actual situation. In our practical experience, is usually set to 500 and is set to 0.7.